OpenStudy (anonymous):

The product of 3 different integers is equal to 7^3. What is the sum?

OpenStudy (rizags):

do you want the answer or an explanation or both

OpenStudy (anonymous):

So far I have done this: x * (x+1) * (x+2) = 7^3 (x^2+x) * (x+2) = 7^3 x^3 + 3x^2 + 2x = 7^3 However when I try to solve for x I get an irrational number. I want both an answer and the explanation.

OpenStudy (rizags):

no good. you're making it way too complicated

OpenStudy (rizags):

first, i want you to evaluate 7^3 and give me all factors in ascending order

OpenStudy (rizags):

btw what grade and what class is this for?

OpenStudy (anonymous):

7^3 = 343 All the factors of 343 are: 1,7,49,343 I am in 10th grade and learning algebra 1 and algebra 2.

OpenStudy (rizags):

ok cool, me too. anyway, since it is identified that those numbers are DIFFERENT, that means that they must all be factors of 343, therefore, i want you to look at those factors and pick out the 3 that work!

OpenStudy (anonymous):

1*7*49=343 So the sum of these factors is 57.

OpenStudy (rizags):

yes! medal plzz and fan